the lenght of a rectangle is three times the width. what is the width if the area is 147 square feet?
Answer by jringelman(2)

start with 2 equations
l=3w
lw=147
use first equation to substitute "3w" for "l" in 2nd equation
(3w)w=147
3w^2=147
w^2=49
w=7 (this is the answer)
verify
l=3w=3*7=21
lw=7*21=147